OpenStudy (anonymous):

Find the number of possible positive and negative real zeros for the function (2x^3)-(x^2)-+5x+2=P(x)

OpenStudy (cp9454):

its -5x or +5x

OpenStudy (cp9454):

i am actually asking about your question its not the answer btw

OpenStudy (anonymous):

THANKS! immmm how'd you get that??

OpenStudy (cp9454):

because u have used both the signs

OpenStudy (anonymous):

oh sorry it's +5x

OpenStudy (cp9454):

first find a root by hit and trial method.

OpenStudy (aum):

Apply Descarte's Rule of Signs.

OpenStudy (goformit100):

I am sorry I didn't understand.... can you please explain it again Sir/Ma'am

OpenStudy (anonymous):

This is literally the question it gives me: Find the number of possible positive and negative real zeros for the function (2x^3)-(x^2)+5x+2=P(x) And I'm a girl.... >:(

OpenStudy (aum):

P(x) = 2x^3 - x^2 + 5x + 2 Descartes's Rule of Signs: How many times does the coefficient change signs? +2 to -1; -1 to +5. Twice. That means there are either 2 positive real roots or zero positive real roots (always goes down in steps of 2).

OpenStudy (anonymous):

okay what about the negative roots??

OpenStudy (anonymous):

i mean zeroes

OpenStudy (aum):

P(-x) = 2(-x)^3 - (-x)^2 + 5(-x) + 2 p(-x) = -2x^3 - x^2 - 5x + 2 How many sign changes? Just once. Therefore, there is definitely one negative real root. Possibilities: 1 negative real root, 2 positive real root, no complex roots OR 1 negative real root, 0 positive real root, 2 complex roots.

OpenStudy (aum):

"roots" "zeros" mean the same thing.

OpenStudy (anonymous):

thanks so much!

OpenStudy (aum):

You are welcome. Since it was a third degree polynomial there will be a total of 3 roots as shown in the answer above.

OpenStudy (goformit100):

Thanks so much from My Side too